OpenStudy (anonymous):

An aircraft has a mass of 2.80 x 105 kg. At a certain instant during its landing, its speed is 51.8 m/s. If the braking force is 3.70 x 105 N, what is the speed of the airplane 13.9 s later?

OpenStudy (anonymous):

This is a motion problem that is best done with a picture:  Solve with \[A = \frac {\Delta v}{\Delta t} = \frac {v_f -v_i}{t_f - t_i} -> v_f = v_i + A(t_f - t_i)\]